What is the total magnification for each lens setting on a microscope with 15x oculars and 4x, 10x, 45x. and 97x objective lenses
Answer #1

The total magnification for each lens setting on a microscope with 15x oculars are-

MT = 15*4=60x

MT = 15*10=150×

MT = 15*45=675x

MT =15*97= 1455x